# Learning an Ensemble Token from Task-driven Priors in Facial Analysis

### v2

@@ 4.2 Datasets
We utilized the UTKFace [46] for age estimation, the AffectNet-HQ [23] for face emotion recognition, the WFLW [43] for face landmark detection, and the CFP-dataset [31] for face recognition. The experimental evaluation utilized four distinct facial analysis datasets: UTKFace, a large-scale repository containing over 20,000 diverse facial images annotated for age, gender, and ethnicity, exhibiting substantial intra-dataset variability; AffectNet, a comprehensive dataset of over one million internet-sourced facial images, with approximately 440,000 manually annotated for seven discrete facial expressions and valence-arousal intensity, designed to facilitate research in automated facial expression recognition; WFLW, a dataset comprising 10,000 in-the-wild facial images with 98 manually annotated landmarks and rich attribute annotations, enabling comprehensive algorithm analysis under significant variations in expression, pose, and occlusion within a unified evaluation framework; and CFP-dataset, consisting of 500 individuals, balanced for gender and aiming for racial diversity, with 10 frontal and 4 profile facial images per individual, rigorously filtered for pose accuracy using human annotation and automated detection. For the scope of this paper, only frontal facial images were considered.
  CFP-dataset
Structure Accuracy $\uparrow$ F1-score $\uparrow$ intra-C variance $\downarrow$ TAR@FAR=0.1% $\uparrow$
M (baseline) 0.8813 0.8504 0.6310 0.3678
M to M 0.8960 (+1.47%) 0.8613 (+1.09%) 0.3186 (-0.312) 0.4938 (+12.60%)
C to M 0.8866 (+0.53%) 0.8602 (+0.98%) 0.3874 (-0.243) 0.4959 (+12.81%)
S to M 0.8960 (+1.47%) 0.8626 (+1.22%) 0.3158 (-0.315) 0.5248 (+15.70%)
C (baseline) 0.9320 0.9136 0.1714 0.7645
M to C 0.9413 (+0.93%) 0.9181 (+0.45%) 0.1632 (-0.008) 0.7893 (+2.48%)
C to C 0.9440 (+1.20%) 0.9243 (+1.07%) 0.1668 (-0.004) 0.7934 (+2.89%)
S to C 0.9466 (+1.46%) 0.9271 (+1.35%) 0.1738 (+0.002) 0.7686 (+0.41%)
S (baseline) 0.8920 0.8584 0.4580 0.6488
M to S 0.8906 (-0.14%) 0.8509 (-0.75%) 0.3493 (-0.108) 0.6921 (+4.33%)
C to S 0.8933 (+0.13%) 0.8644 (+0.60%) 0.4518 (-0.006) 0.6302 (-1.86%)
S to S 0.8973 (+0.53%) 0.8576 (-0.08%) 0.5966 (+0.138) 0.6507 (+0.82%)
Table 1: Evaluation table on CFP-dataset.
@@
Figure 6: Visualization of embeddings in CFP-dataset. Each color delineates a unique category within individual person identity. The elliptical contours visually depict the intra-class variance for each respective class. As corroborated by  1, the intra-class variance and the distinction of classes is significantly improved in the  6(b) compared to the  6(a)
@@ 4.3 Pre-training the Prior Models
For illustrative examples as experimental table, ”M (baseline)” denotes the architectural structure pre-trained on MobileOne, while ”C to M” denotes a configuration wherein ConvNeXt serves as the pre-trained model for the N-1 prior branches, and MobileOne is designated as the canonical branch for knowledge token learning. For visualizing the embeddings, we employed the t-SNE [35] to reduce the N-dimensional embeddings to two-dimensional embeddings in  7 and  6.
4.4 Evaluation per Datasets
  WFLW
@@ 4.4 Evaluation per Datasets
Table 2: Evaluation table on WFLW dataset. Red text indicates performance improvements, while blue text indicates performance degradation.
To assess face landmark detection performance on the WFLW dataset, normalized mean error (NME) and mean absolute error (MAE) on 98 landmark points were adopted as evaluation metrics. The NME was normalized by the inter-ocular distance, defined as the Euclidean distance between the lateral canthi of the left and right eyes. The MAE was scaled relative to the input size of 224. Although significant performance enhancements were not universally demonstrated in  2, all metrics exhibited a statistically significant improvement with the KT-Adapter methodology, with the exception of NME of ”S to M” with +0.004. Notably, the adoption of CNN-based task priors consistently yielded more pronounced performance compared to ViT-based priors.
(a) M (baseline)
(b) C to M (ours)
Figure 7: Visualization of embeddings (token) in UTK-face dataset. The color of each data point corresponds to the age represented in the right color-bar. Notably, data points that were distinct within the red ellipse in  7(a) are not observed in the ”C to M” data points in  7(b). This observation suggests that the KT-Adapter not only assimilate the corner-case scenarios but also represent the feature clustering with adaptable transforms.
  UTK-face
@@
For the evaluation of age estimation performance, MAE and cumulative score (CS) metrics were employed on the UTK-face dataset. The MAE was computed as the absolute deviation between predicted and ground truth ages, spanning the range of 1 to 116 years. The CS metric was determined as the percentage of predictions falling within $\pm$2 years, a tolerance level of 5. Consistent with observations in landmark detection, MAE exhibited marginal performance enhancements, lacking substantial improvements in  1. However, the CS metric demonstrated statistically significant performance gains. Notably, the adaptation of task priors among high-capacity models within the ConvNeXt architecture and Swin Transformer yielded superior performance improvements, ”C to C” with +0.78% and ”S to C” +0.98%, compared to alternative architectures, which achieves substantial performance gains relative to KT-Adapter’s minimal computational overhead.
  Affect-Net-HQ
@@
M (baseline) 0.7261 0.7023
M to M 0.7309 (+0.48%) 0.7052 (+0.29%)
C to M 0.7299 (+0.38%) 0.7040 (+0.17%)
S to M 0.7323 (+0.62%) 0.7063 (+0.40%)
C (baseline) 0.7467 0.7233
M to C 0.7434 (-0.33%) 0.7204 (-0.29%)
C to C 0.7465 (-0.02%) 0.7235 (+0.02%)
S to C 0.7472 (+0.05%) 0.7245 (+0.12%)
S (baseline) 0.7390 0.7176
M to S 0.7513 (+1.23%) 0.7276 (+1.00%)
C to S 0.7520 (+1.30%) 0.7277 (+1.01%)
S to S 0.7501 (+1.11%) 0.7263 (+0.87%)
@@
For the classification task of face emotion recognition, we employed standard evaluation metrics, accuracy and F1-score. As shown in 4, all methodologies, with the exception of ConvNeXt, exhibited performance enhancements. We posit that this phenomenon is attributable to the inherent high capacity of ”C” for the emotion recognition task relative to the other task priors. Notably, despite ”S” demonstrating significantly lower accuracy compared to ”C” (0.7390 and 0.7176 vs. 0.7467 and 0.7233), the application of KT-Adapter in the ”C to S” outperformed with superior performance (0.7520 and 0.7277). This outcome is particularly compelling, as the observed performance gains in the ”S to C”, where the task priors are reversed to ”C to S”, suggest a synergistic relationship between ConvNeXt and Swin for face emotion recognition.
@@
Table 5: Ablation study on parameter N, the number of encoders. We reported the results using the best-performing encoder among those sharing the same architecture type and encoder count. For example, in the face recognition task, the best performance was achieved when using encoders on age estimation and face landmark detection. We observed that increasing the number of prior encoders generally led to a linear improvement in performance.
@@ 4.5 Ablation Studies
An ablation study was conducted to assess the effects of the number of encoders N in  5, the use of drop mask ”DM”, canonical mask ”CM”, and the positional encoding ”PE” in  6. Evaluation metrics included accuracy and F1-score for facial emotion recognition, and CS@5 and MAE for age estimation. The use of the drop mask alone yielded improvements compared to the baseline model. Applying the canonical mask alone also yielded performance gains. Notably, a synergistic effect was observed when both the drop mask and canonical mask were employed simultaneously. As discussed in Section  3.4, this performance improvement indicates that structural misalignment between the training and inference phases plays a significant role in KT-Adapter. Interestingly, incorporating positional encoding into the prior tokens had only a minimal impact on performance. Considering the computational efficiency benefits of drop mask and canonical mask in training, we exclude positional encoding to reduce computational overhead.
@@ 5 Conclusion
The proposed KT-Adapter framework exhibited performance gains across tasks and evaluation metrics, while maintaining a demonstrably minimal computational overhead, with particularly notable improvements observed at the feature representation level. However, a salient limitation inherent in knowledge token methodologies is their susceptibility to overfitting datasets. The training phase, which relies on task priors and presents each encoder with canonical input, potentially impedes the learning of multi-modal or multi-input modalities. We posit that this research serves as a cornerstone towards the effective exploitation of task priors and encourage future investigations to broaden its applicability beyond the domain of facial analysis, extending to more generalized domains.

### v3

@@ 4.2 Datasets
We utilized the UTK-face [zhang2017ageprogressionregressionconditionaladversarial] for age estimation, the AffectNet-HQ [8013713] for face emotion recognition, the WFLW [wayne2018lab] for face landmark detection, and the CFP-dataset [7477558] for face recognition. The experimental evaluation utilized four distinct facial analysis datasets: UTK-face, a large-scale repository containing over 20,000 diverse facial images annotated for age, gender, and ethnicity, exhibiting substantial intra-dataset variability; AffectNet, a comprehensive dataset of over one million internet-sourced facial images, with approximately 440,000 manually annotated for seven discrete facial expressions and valence-arousal intensity, designed to facilitate research in automated facial expression recognition; WFLW, a dataset comprising 10,000 in-the-wild facial images with 98 manually annotated landmarks and rich attribute annotations, enabling comprehensive algorithm analysis under significant variations in expression, pose, and occlusion within a unified evaluation framework; and CFP-dataset, consisting of 500 individuals, balanced for gender and aiming for racial diversity, with 10 frontal and 4 profile facial images per individual, rigorously filtered for pose accuracy using human annotation and automated detection. For the scope of this paper, only frontal facial images were considered.
@@
Figure 6: Visualization of embeddings in CFP-dataset. Each color delineates a unique category within individual person identity. The elliptical contours visually depict the intra-class variance for each respective class. As corroborated by Tab. 3, the intra-class variance and the distinction of classes is significantly improved in the Fig. 6(b) compared to the Fig. 6(a)
(a) M (baseline)
(b) C to M (ours)
Figure 7: Visualization of embeddings (token) in UTK-face dataset. The color of each data point corresponds to the age represented in the right color-bar. Notably, data points that were distinct within the red ellipse in Fig. 7(a) are not observed in the ”C to M” data points in Fig. 7(b). This observation suggests that the KT-Adapter not only assimilate the corner-case scenarios but also represent the feature clustering with adaptable transforms.
@@ 4.3 Pre-training the Prior Models
For illustrative examples as experimental table, ”M (baseline)” denotes the architectural structure pre-trained on MobileOne, while ”C to M” denotes a configuration wherein ConvNeXt serves as the pre-trained model for the N1N-1 prior branches, and MobileOne is designated as the canonical branch for knowledge token learning. For visualizing the embeddings, we employed the t-SNE [van2008visualizing] to reduce the N-dimensional embeddings to two-dimensional embeddings in Fig. 7 and Fig. 6.
4.4 Evaluations on target task
As shown in Tab. 1, in order to ensure a fair comparison, all comparative models were trained with encoders frozen while only the decoders were fine-tuned. All encoder architectures were kept identical to that of the baseline. The results indicate that both Model Ensemble [lakshminarayanan2017deepensembles] and Model Soup [pmlr-v162-wortsman22a] exhibit instability when the pre-processing procedures and target tasks are not aligned, especially on landmark detection and age estimation. In contrast, our method demonstrates remarkable robustness, maintaining strong performance even when trained across datasets and tasks.
@@ 4.4 Evaluations on target task
Table 2: Evaluation table on WFLW dataset.
  Structure intra-C variance $\downarrow$ TAR@FAR=0.1% $\uparrow$
M (baseline) 0.6310 0.3678
M to M 0.3186 (-0.312) 0.4938 (+12.60%)
C to M 0.3874 (-0.243) 0.4959 (+12.81%)
S to M 0.3158 (-0.315) 0.5248 (+15.70%)
C (baseline) 0.1714 0.7645
M to C 0.1632 (-0.008) 0.7893 (+2.48%)
C to C 0.1668 (-0.004) 0.7934 (+2.89%)
S to C 0.1738 (+0.002) 0.7686 (+0.41%)
S (baseline) 0.4580 0.6488
M to S 0.3493 (-0.108) 0.6921 (+4.33%)
C to S 0.4518 (-0.006) 0.6302 (-1.86%)
S to S 0.5966 (+0.138) 0.6507 (+0.82%)
Table 3: Evaluation table on CFP-dataset.
To assess face landmark detection performance on the WFLW dataset, normalized mean error (NME) and mean absolute error (MAE) on 98 landmark points were adopted as evaluation metrics. The NME was normalized by the inter-ocular distance, defined as the Euclidean distance between the lateral canthi of the left and right eyes. The MAE was scaled relative to the input size of 224. Although significant performance enhancements were not universally demonstrated in Tab. 2, all metrics exhibited a statistically significant improvement with the KT-Adapter methodology, with the exception of NME of ”S to M” with +0.004. Notably, the adoption of CNN-based task priors consistently yielded more pronounced performance compared to ViT-based priors.
@@
M (baseline) 72.61 70.23
M to M 73.09 (+0.48%) 70.52 (+0.29%)
C to M 72.99 (+0.38%) 70.40 (+0.17%)
S to M 73.23 (+0.62%) 70.63 (+0.40%)
C (baseline) 74.67 72.33
M to C 74.34 (-0.33%) 72.04 (-0.29%)
C to C 74.65 (-0.02%) 72.35 (+0.02%)
S to C 74.72 (+0.05%) 72.45 (+0.12%)
S (baseline) 73.90 71.76
M to S 75.13 (+1.23%) 72.76 (+1.00%)
C to S 75.20 (+1.30%) 72.77 (+1.01%)
S to S 75.01 (+1.11%) 72.63 (+0.87%)
@@
For the evaluation of age estimation performance, MAE and cumulative score (CS) metrics were employed on the UTK-face dataset. The MAE was computed as the absolute deviation between predicted and ground truth ages, spanning the range of 1 to 116 years. The CS metric was determined as the percentage of predictions falling within $\pm$2 years, a tolerance level of 5. Consistent with observations in landmark detection, MAE exhibited marginal performance enhancements, lacking substantial improvements in Tab. 3. However, the CS metric demonstrated statistically significant performance gains. Notably, the adaptation of task priors among high-capacity models within the ConvNeXt architecture and Swin Transformer yielded superior performance improvements, ”C to C” with +0.78% and ”S to C” +0.98%, compared to alternative architectures, which achieves substantial performance gains relative to KT-Adapter’s minimal computational overhead.
@@
Table 6: Ablation study on parameter NN, the number of encoders. We reported the results using the best-performing encoder among those sharing the same architecture type and encoder count. For example, in the face recognition task, the best performance was achieved when using encoders on age estimation and face landmark detection. We observed that increasing the number of prior encoders generally led to a linear improvement in performance.
@@
For the classification task of face emotion recognition, we employed standard evaluation metrics, accuracy and F1-score. As shown in Tab. 5, all methodologies, with the exception of ConvNeXt, exhibited performance enhancements. We posit that this phenomenon is attributable to the inherent high capacity of ”C” for the emotion recognition task relative to the other task priors. Notably, despite ”S” demonstrating significantly lower accuracy compared to ”C” (0.7390 and 0.7176 vs. 0.7467 and 0.7233), the application of KT-Adapter in the ”C to S” outperformed with superior performance (0.7520 and 0.7277). This outcome is particularly compelling, as the observed performance gains in the ”S to C”, where the task priors are reversed to ”C to S”, suggest a synergistic relationship between ConvNeXt and Swin for face emotion recognition.
@@ 4.5 Ablation Studies
An ablation study was conducted to assess the effects of the number of encoders N in Tab. 6, the use of drop mask ”DM”, canonical mask ”CM”, and the positional encoding ”PE” in Tab. 7. Evaluation metrics included accuracy and F1-score for facial emotion recognition, and CS@5 and MAE for age estimation. The use of the drop mask alone yielded improvements compared to the baseline model. Applying the canonical mask alone also yielded performance gains. Notably, a synergistic effect was observed when both the drop mask and canonical mask were employed simultaneously. As discussed in Sec. 3.4, this performance improvement indicates that structural misalignment between the training and inference phases plays a significant role in KT-Adapter. Interestingly, incorporating positional encoding into the prior tokens had only a minimal impact on performance. Considering the computational efficiency benefits of drop mask and canonical mask in training, we exclude positional encoding to reduce computational overhead.
@@ 5 Conclusion
The KT-Adapter exhibited performance gains across tasks and evaluation metrics, while maintaining a demonstrably minimal computational overhead, with particularly notable improvements observed at the feature representation level. However, a salient limitation inherent in knowledge token methodologies is their susceptibility to overfitting datasets. The training phase, which relies on task priors and presents each encoder with canonical input, potentially impedes the learning of multi-modal or multi-input modalities. We posit that this research serves as a cornerstone towards the effective exploitation of task priors and encourage future investigations to broaden its applicability beyond the domain of facial analysis, extending to more generalized domains.